OpenStudy (anonymous):

Guys, can you give me spare time helping me answer this Linear Equations with 2 Variables By Add a Subt 1a. x-3y=2 2x+3y=13 2b. 2/x+3/y=7 6/x-6/y=1

OpenStudy (anonymous):

1a. x - 3y = 2 2x+ 3y=13 3x = 15 3x/3 = 15/3 x = 5

OpenStudy (anonymous):

Any questions?

OpenStudy (anonymous):

I have to go.

OpenStudy (amistre64):

your step thru is a bit confusing

OpenStudy (amistre64):

i beleive that there is a "method" that is being asked to apply .... substitution perhaps?

OpenStudy (anonymous):

elimination

OpenStudy (amistre64):

or is that: by addition or substitution?

OpenStudy (amistre64):

elimination will do fine

OpenStudy (amistre64):

can you tell me what the easiest elimination in part1 is?

OpenStudy (amistre64):

2x+2x is not 0x

OpenStudy (amistre64):

x - 3y = 2 <-- * -2 2x + 3y = 13 -2x +6y = -4 2x + 3y = 13 -------------- 0x + 9y = 9

OpenStudy (anonymous):

ill show you how i did it x-3y=2 2x+3y=13 x - 2 (x-3y=2) x- 1(2x+3y=13) 2x-6y= 4 2x+(-)3=(-)13 9y 9 9 9 y=1

OpenStudy (anonymous):

so the next part im confused

OpenStudy (amistre64):

x-3y=2 2x+3y=13 x - 2 (x-3y=2) x- 1(2x+3y=13) 2x-6y= 4 (-)2x+(-)3=(-)13 <--- you forgot a negative .... 9y 9 9 9 y=1 , now this makes sense

OpenStudy (anonymous):

ok thanks for that, can you help me whats next?

OpenStudy (anonymous):

to find x

OpenStudy (amistre64):

notice that the ys already eliminate themsleves, so just add em up vvv x - 3y = 2 2x + 3y = 13

OpenStudy (amistre64):

OR, since you already know that y=1 x - 3y = 2 x - 3(1) = 2

OpenStudy (anonymous):

no i think after that its this x 1 (x-3y=2) x 1 (2x+3y=13) how to find x here?

OpenStudy (anonymous):

can you help me out with number 2?